On a rainy day, a girl broke up with her boyfriend after being together for eight years. They decided to separate at the place where everything about them began, at the same time. The boy is due north crying and running at a rate of 5 ft/sec and the girl is walking due east at rate of 1 ft/sec thinking if she made the right decision. How fast are they separating from each other 5 seconds after they started moving to a new life without each other?


1
Expert's answer
2021-07-25T09:17:21-0400

"v=\frac{v_x x+v_y y}{d}=\frac{1\cdot 5+5\cdot 25}{\sqrt{5^2+25^2}}=\sqrt{26}\approx 5.1~\frac {ft}{s}."